OpenStudy (anonymous):

-7d<_ 4d-3<_-1 help pls!!

OpenStudy (anonymous):

Pls help me i beg you

hero (hero):

\[-7d \le 4d - 3 \le - 1\]

hero (hero):

Is that it?

OpenStudy (anonymous):

Yes it is

hero (hero):

Split it into two , then solve.

hero (hero):

@captainmico88, did you do it yet bro?

OpenStudy (anonymous):

I cant

OpenStudy (anonymous):

I dont get this one bro

OpenStudy (anonymous):

Help men

hero (hero):

\[4d - 3 \ge-7d\] \[4d - 3 \le -1\]

hero (hero):

@captainmico88 Solve each separately

hero (hero):

\[4d - 3 \ge-7d\] \[4d - 3 \le -1\]

hero (hero):

@captainmico88, did you do it yet bro?

OpenStudy (anonymous):

It get >_ 3.6 int the first